### v1

@@ 3 CGP-CDE: Causal Gaussian Process Conditional Density Estimators
We propose a method that scales up Bayesian model selection based causal discovery by using the insights of Zheng et al. (2018) (section 2.5). We first describe the model that we use, then we show how the model can continuously parameterise the space of graphs. This is followed by a loss function that maximises the posterior probability and ensures that the final graph is a DAG.
@@ 3.1 The GP-CDE Model
The Bayesian model selection approach to causal structure learning described in section 2.4 requires defining distributions and priors for a causal model. To take advantage of the Bayesian framework, we wish to use a model for each $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ that is more flexible than previous attempts. For this, we use a version of the Gaussian process conditional density estimator (GP-CDE) model (Dutordoir et al., 2018).
@@
where $\mathbf{W}\in\mathbb{R}^{N\times D}$ are the latent variables, $\boldsymbol{\phi}\in\mathbb{R}^{D}$ are the likelihood variances, $\boldsymbol{\Lambda}_{i}$ are the kernel hyperparameters, and $\mathcal{M}_{\mathcal{G}}$ denotes the causal model with DAG $\mathcal{G}$. Each variable $\mathbf{x}_{i}$ is modelled as the output of a function $\mathbf{f}_{i}$, with the parents of the variable in $\mathcal{G}$, $\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)}$, and a latent variable $\mathbf{w}_{i}$ being the inputs. The inclusion of $\mathbf{w}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ allows for each variable to be described with non-Gaussian and heteroscedastic noise, greatly increasing the expressivity of the model (Dutordoir et al., 2018). The term $p(\mathbf{f}_{i}|\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)},\boldsymbol{\Lambda}_{i},w_{ni})$ is the Gaussian process prior (Rasmussen, 2003), and equal to
@@ 3.2 Continuous Relaxation: The CGP-CDE Model
The key insight to continuously parameterise graphs in the above model (not DAGs), is that the hyperparameters control dependence between variables (Williams and Rasmussen, 1995). In our Gaussian process prior for a variable $X_{i}$, we use kernels with different hyperparameters $\Lambda_{ij}$ for each dimension $j$). In general, kernels of this form factorise as $K_{\boldsymbol{\Lambda}_{i}}\mathopen{}\mathclose{{}\left((\mathbf{X},\mathbf{w}_{i}),(\mathbf{X},\mathbf{w}_{i})^{\prime}}\right)=K_{\Lambda_{ii}}\mathopen{}\mathclose{{}\left(\mathbf{w}_{i},\mathbf{w}_{i}^{\prime}}\right)\prod_{j\neq i}K_{\Lambda_{ij}}\mathopen{}\mathclose{{}\left(\mathbf{x}_{j},\mathbf{x}_{j}^{\prime}}\right)$. Certain hyperparameters of the kernels can then be used to control the variability of the function with respect to specific inputs. We denote these hyperparameters as $\boldsymbol{\theta}$ with $\boldsymbol{\theta}\subset\boldsymbol{\Lambda}$ and all other hyperparameters as $\boldsymbol{\sigma}=\boldsymbol{\Lambda}\setminus\boldsymbol{\theta}$. Specifically
@@
Hence, if the kernel hyperparameter value for $\theta_{ij}$ is near zero, the function for variable $X_{i}$ is constant with respect to variable $X_{j}$. Thus we can construct a corresponding adjacency as
@@
As a value of $\theta_{ij}=0$ means that changing $X_{j}$ does not change $X_{i}$ this corresponds to an absence of an edge from $X_{j}$ to $X_{i}$ in $\mathbf{A}$.
@@
where $\mathrm{PA}_{\mathcal{G}^{\mathbf{A}}}(i)$ denotes the parents of $X_{i}$ in $\mathcal{G}^{\mathbf{A}}$ (the graph implied by the adjacency $\mathbf{A}$), and $\mathbf{X}_{\neg i}$ denotes all variables except $X_{i}$. Thus, for a variable $X_{i}$, all the other variables are inputs, and the values of the hyperparameters $\boldsymbol{\theta}_{i}$ control the dependence of the rest of the variables $\mathbf{X}_{\neg i}$ on $X_{i}$, and hence define the parents of $X_{i}$. The hyperparameters in the above model thus continuously parameterise the space of graphs. The exact kernel we use is in section B.1.
@@ 3.3 Priors on Graphs
As discussed in Eggeling et al. (2019), not using a prior on graphs leads to a higher weight on denser graphs due to the larger number of dense graphs versus sparse graphs. Hence, to even out this effect, we place a prior that prefers sparser graphs. By introducing priors on the hyperparameters $\boldsymbol{\theta}$ in our model, we can effectively encode priors on the graph in eq. 9. We thus place a Gamma prior with parameters chosen to prefer small values of $\boldsymbol{\theta}$, $P(\boldsymbol{\theta})=\text{Gamma}(\eta,\beta)$. Note that our prior is symmetric between graphs within the same MEC. That is, we do not impose a preference over the causal direction between variables, only over the number of edges in the graph. Other priors on graphs could be considered (Eggeling et al., 2019), but we leave that for future work.
@@ 3.4 Score
As stated in section 2.4, our decision rule is to pick the causal graph with the highest posterior probability. For this, we need to calculate the log marginal likelihood for every variable in our model (eq. 3). To find the log marginal likelihood, we must integrate over the priors of $\mathbf{f}$ and $\mathbf{W}$, $\log p(\mathbf{x}_{i}|\mathbf{X}_{\mathrm{PA}_{\mathcal{G}^{\mathbf{A}}}(i)},\boldsymbol{\sigma}_{i},\boldsymbol{\theta}_{i},\phi_{i})$ is given as
@@
Calculating this marginal likelihood is analytically intractable due to the non-linear dependence of $\mathbf{f}_{i}$ on the latent term $\mathbf{w}_{i}$. We thus use variational inference, with variational posteriors $q$, to optimise a tractable lower bound to the marginal likelihood (details and derivation as well as how we handle the rest of the hyperparameters are in section B.2). We denote the full lower bound for all variables as $\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi})$. The optimisation problem is
@@
We can use the function in eq. 5 to convert the condition that $\mathcal{G}^{\mathbf{A}_{\boldsymbol{\theta}}}\in\text{DAGs}$ to $h(\mathbf{A}_{\boldsymbol{\theta}})=0$. The constrained optimisation problem can now be transformed into a sequence of unconstrained optimisation problems, solved using the augmented Lagrange (section B.3) (Zheng et al., 2018). The final score is
$\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi})\!+\!\log p(\boldsymbol{\theta})\!-\!\alpha_{t}|h(\mathbf{A}_{\boldsymbol{\theta}})|^{2}\!-\!\frac{\rho_{t}}{2}h(\mathbf{A}_{\boldsymbol{\theta}}).$ (11)
Structure learning has now been converted into a single differentiable optimisation problem.
@@ 3.5 Computational Cost
The cost of computing the loss for $N$ samples in a Gaussian Process is usually $\mathcal{O}(N^{3})$. We use the uncollapsed inducing point formulation that introduces $M<N$ inducing points allowing for mini-batching (section B.2)(Hensman et al., 2013). This changes the cost to $\mathcal{O}(M^{3})$. As we have $D$ variables, the total cost is $\mathcal{O}(DM^{3})$. The cost of computing the acyclicity regulariser is $\mathcal{O}(D^{3})$ for $D$ variables. Hence, the cost is dominated by $M$ or $D$, whichever is bigger. The choice of the number of inducing points depends on the problem, with more inducing points leading to better approximations (Bauer et al., 2016). However, it is possible to use the lower bound $\mathcal{L}_{\mathrm{ELBO}}$ to find the optimal number of inducing points (Burt et al., 2020). The cost of Gaussian processes, and Bayesian methods in general, can be higher than their non-Bayesian counterparts. Despite this, Bayesian model selection based causal discovery offers other advantages such as the use of more flexible model classes.
Figure 2: Box plots of metrics for (a) three variable synthetic data (b) Erdos-Renyi (ER) graphs (c) Scale-Free (SF) graphs (d) Syntren data. The plots show the distribution of the metrics across the different datasets for each experiment. The centre line is the median, the black dot is the mean and the white dots represent outliers. Lower is better for SHD and SID, and higher is better for F1.
@@ 4 Related Work
Certain approaches only use information about independences between variables to discover the causal structure, recovering only an MEC, and not the whole DAG (Spirtes et al., 2000; 1995; Chickering, 2002; Huang et al., 2018). Identifying a unique DAG requires additional assumptions, typically involving restrictions on model classes or access to interventional data (Lippe et al., 2021; Ke et al., 2022; Brouillard et al., 2020; Ke et al., 2020). Since we assume access only to observational data, we focus on the former. Continuous optimisation methods cast causal discovery as a single unconstrained optimisation problem (Zheng et al., 2018). This involves either introducing a regulariser to promote learning DAGs (Zheng et al., 2018; Lachapelle et al., 2019; Yu et al., 2019) or directly parameterising the space of DAGs (Yu et al., 2021; Charpentier et al., 2021). NOTEARS (Zheng et al., 2018) uses a regulariser for learning linear causal models. Follow-up work, such as GraN-DAG (Lachapelle et al., 2019) and DAG-GNN (Yu et al., 2019), extends this approach to neural networks and graph neural networks. These methods typically restrict models to additive noise (ANM) to recover a DAG, which can limit performance when assumptions are not met (section 2.3).
@@
The Bayesian framework has previously been used for causal discovery, focusing on efficient sampling from the posterior over causal graphs. Geiger and Heckerman (2002) use Bayesian linear Gaussian models but construct their model such that graphs in the same Markov equivalence class receive the same score. Similarly, Cundy et al. (2021) also only consider posteriors over linear Gaussian models. DiBS (Lorch et al., 2021) assumes Bayesian ANMs and parametrises the space of graphs using latent variables. Their method is able to output a fixed number of samples from the posterior over causal graphs. Annadani et al. (2023) also use sampling methods to sample from the posterior over ANM causal models but use a permutation based parametrisation of the graph (Charpentier et al., 2021; Yu et al., 2021).
In contrast, our work uses Bayesian model selection to distinguish causal structures within an MEC that can allow for fewer restrictions than likelihood-based methods (Stegle et al., 2010; Dhir et al., 2024). While previous studies demonstrated this with two variables, we show its feasibility and advantages with multiple variables.
@@ 5 Experiments
We compare our approach to various baselines that perform multivariate causal discovery to recover a DAG. We test our method on synthetic data generated from our model (section 5.1), and data not generated from our model (section 5.2). Then we compare our models to the baselines in semi-synthetic and real data cases (section 5.3). See section B.6 for full details of the settings used for the CGP-CDE implementation.
It was recently found that causal discovery methods can be sensitive to variance information that is an artefact of synthetic causal datasets or measurement scales in real world datasets (Reisach et al., 2021). Following suggestions in Reisach et al. (2021), we standardise every variable before using our method and the baselines. We note that previous works did not standardise all the datasets, explaining the difference in results from our work (see appendix D).
Baselines: We compare against non-Bayesian counterparts that use similar acyclicity regularisers, such as NOTEARS (Zheng et al., 2018), GraN-DAG, and GraN-DAG++ (Lachapelle et al., 2019), providing a direct comparison of our method’s advantage. NOTEARS assumes linearity, GraN-DAG assumes ANM, and GraN-DAG++ is not identifiable. We also compare against order based methods that first estimate a topological order of the graph and then prune to get the final graph - SCORE (Rolland et al., 2022), NOGAM (Montagna et al., 2023), and CAM (Bühlmann et al., 2014). These all assume ANM. Details are in appendix F.
Metrics: We use the following metrics to compare the methods. SHD: The structural Hamming distance is the Hamming distance between the predicted adjacency and the ground truth adjacency matrix. SID: The structural interventional distance counts the number of interventional distributions that are incorrect if the predicted graph is used to form the parent adjustment set instead of the ground truth graph (Peters and Bühlmann, 2015). F1: The F1 score is the harmonic mean of the precision and recall, where an edge is considered the positive class.

### v2

@@ 4 CGP-CDE: Causal Gaussian Process Conditional Density Estimators
Although the posterior can allow for more accurate causal discovery with fewer restrictions, computing and comparing the posterior of every causal structure becomes intractable when the number of variables is increased. We propose a method that scales up Bayesian model selection based causal discovery by using the insights of Zheng et al. (2018) (Section A.2). We first describe the model that we use, then we show how the model can continuously parameterise the space of graphs. This is followed by a loss function that maximises the posterior probability and ensures that the final graph is a DAG.
@@ 4.1 The GP-CDE Model
The Bayesian model selection approach to causal structure learning described in Section 2.4 requires defining distributions and priors for a causal model. To take advantage of the Bayesian framework, we wish to use a model for each $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ that is more flexible than previous attempts. For this, we use a version of the Gaussian process conditional density estimator (GP-CDE) model (Dutordoir et al., 2018). Here, $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ is a class of nonparametric functions and the prior over the conditionals is a Gaussian process prior.
@@
where $\mathbf{W}\in\mathbb{R}^{N\times D}$ are the latent variables, $\boldsymbol{\phi}\in\mathbb{R}^{D}$ are the likelihood variances, $\boldsymbol{\Lambda}_{i}$ are the kernel hyperparameters, and $\mathcal{M}_{\mathcal{G}}$ denotes the causal model with DAG $\mathcal{G}$. Each variable $\mathbf{x}_{i}$ is modelled as the output of a function $\mathbf{f}_{i}$, with the parents of the variable in $\mathcal{G}$, $\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)}$, and a latent variable $\mathbf{w}_{i}$ being the inputs. The inclusion of $\mathbf{w}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ allows for each variable to be described with non-Gaussian and heteroscedastic noise, greatly increasing the expressivity of the model (Dutordoir et al., 2018) [^3] Our model can be thought of as a Bayesian conditional variational autoencoder.. The term $p(\mathbf{f}_{i}|\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)},\boldsymbol{\Lambda}_{i},w_{ni})$ is the Gaussian process prior (Rasmussen, 2003), and equal to
@@ 4.2 Continuous Relaxation: The CGP-CDE Model
The key insight to continuously parameterise graphs (not DAGs) in the above model is that the hyperparameters control dependence between variables (Williams & Rasmussen, 1995). In our Gaussian process prior for a variable $X_{i}$, we use kernels with different hyperparameters $\Lambda_{ij}$ for each dimension $j$. In general, kernels of this form factorise as $K_{\boldsymbol{\Lambda}_{i}}\mathopen{}\mathclose{{}\left((\mathbf{X},\mathbf{w}_{i}),(\mathbf{X},\mathbf{w}_{i})^{\prime}}\right)=K_{\Lambda_{ii}}\mathopen{}\mathclose{{}\left(\mathbf{w}_{i},\mathbf{w}_{i}^{\prime}}\right)\prod_{j\neq i}K_{\Lambda_{ij}}\mathopen{}\mathclose{{}\left(\mathbf{x}_{j},\mathbf{x}_{j}^{\prime}}\right)$. Certain hyperparameters of the kernels can then be used to control the variability of the function with respect to specific inputs. We denote these hyperparameters as $\boldsymbol{\theta}$ with $\boldsymbol{\theta}\subset\boldsymbol{\Lambda}$ and all other hyperparameters as $\boldsymbol{\sigma}=\boldsymbol{\Lambda}\setminus\boldsymbol{\theta}$. Specifically
@@
Hence, if the kernel hyperparameter value for $\theta_{ij}$ is near zero, the function for variable $X_{i}$ is constant with respect to variable $X_{j}$. This implies an adjacency matrix
@@
A value of $\theta_{ij}=0$ implies that changing $X_{j}$ does not change $X_{i}$ and results in an absence of the edge $X_{j}\to X_{i}$ in $\mathbf{A}$.
@@
where $\mathrm{PA}_{\mathcal{G}^{\mathbf{A}}}(i)$ denotes the parents of $X_{i}$ in $\mathcal{G}^{\mathbf{A}}$ (the graph implied by the adjacency $\mathbf{A}$), and $\mathbf{X}_{\neg i}$ denotes all variables except $X_{i}$. Thus, for a variable $X_{i}$, all the other variables are inputs, and the values of the hyperparameters $\boldsymbol{\theta}_{i}$ control the dependence of the rest of the variables $\mathbf{X}_{\neg i}$ on $X_{i}$, and hence define the parents of $X_{i}$. The hyperparameters in the above model thus continuously parameterise the space of graphs. The exact kernel we use is in Section C.1.
@@ 4.3 Priors on Graphs
As discussed in Eggeling et al. (2019), the lack of a prior on graphs leads to a higher weight on denser graphs due to the larger number of dense graphs versus sparse graphs. Hence, to even out this effect, we place a prior on graphs. By introducing priors on the hyperparameters $\boldsymbol{\theta}$ in our model, we can effectively encode priors on the graph in Equation 10. We thus place a Gamma prior with parameters chosen to prefer small values of $\boldsymbol{\theta}$, $P(\boldsymbol{\theta})=\text{Gamma}(\eta,\beta)$. Note that our prior is symmetric between graphs within the same MEC. That is, we do not impose a preference over the causal direction between variables, only over the number of edges in the graph. Other priors on graphs could be considered (Eggeling et al., 2019), but we leave that for future work.
@@ 4.4 Score
As stated in Section 2.4, our decision rule is to pick the causal graph with the highest posterior probability. For this, we need to calculate the log marginal likelihood for every variable in our model (Equation 6). To find the log marginal likelihood, we must integrate over the priors of $\mathbf{f}$ and $\mathbf{W}$, $\log p(\mathbf{x}_{i}|\mathbf{X}_{\mathrm{PA}_{\mathcal{G}^{\mathbf{A}}}(i)},\boldsymbol{\sigma}_{i},\boldsymbol{\theta}_{i},\phi_{i})$ is given as
@@
Calculating this marginal likelihood is analytically intractable due to the non-linear dependence of $\mathbf{f}_{i}$ on the latent term $\mathbf{w}_{i}$. We thus use variational inference, with variational posteriors $q$, to optimise a tractable lower bound to the marginal likelihood (details and derivation as well as handling of other hyperparameters are in Section C.2). We denote the full lower bound for all variables as $\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi})$. The optimisation problem is
@@
To enforce the condition that $\mathcal{G}^{\mathbf{A}_{\boldsymbol{\theta}}}\in\text{DAGs}$ we use the spectral acyclicity constraint method (Lee et al., 2019)
$h(\mathbf{A_{\boldsymbol{\theta}}})=|\lambda_{d}(\mathbf{A_{\boldsymbol{\theta}}})|,$ (12)
where $\lambda_{d}(\mathbf{A_{\boldsymbol{\theta}}})$ denotes the largest eigenvalue of $\mathbf{A_{\boldsymbol{\theta}}}$, which has been shown to be more stable and scalable than other acyclic constraints (Nazaret et al., 2024). Following Nazaret et al. (2024), a penalty method is used to allow for continuous optimisation, making the final loss
$\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi})\!+\!\log p(\boldsymbol{\theta})\!-\gamma_{t}h(\mathbf{A}_{\boldsymbol{\theta}}),$ (13)
where the weighting $\gamma_{t}$ is increased by $\rho$ at each epoch. A solution $\mathbf{A}^{*}$ is found when $h(\mathbf{A}^{*})<\tau$ for some $\tau>0$. See Section A.2 for more details.
@@ 4.5 Computational Cost
The cost of computing the loss for $N$ samples in a Gaussian Process is usually $\mathcal{O}(N^{3})$. We use the uncollapsed inducing point formulation that introduces $M<N$ inducing points allowing for mini-batching (Section C.2)(Hensman et al., 2013). This changes the cost to $\mathcal{O}(M^{3})$. As we have $D$ variables, the total cost is $\mathcal{O}(DM^{3})$. The cost of computing the acyclicity regulariser is $\mathcal{O}(D^{2})$. Hence, the cost is dominated by $M$. The choice of $M$ depends on the problem, with more inducing points leading to better approximations (Bauer et al., 2016). However, it is possible to use $\mathcal{L}_{\mathrm{ELBO}}$ to find the optimal number of inducing points (Burt et al., 2020). Gaussian processes, and Bayesian methods in general, can be more computationally expensive than their non-Bayesian counterparts. However, Bayesian model selection-based causal discovery provides key advantages, including the ability to use more flexible model classes.
Figure 1: Boxplot of metrics for the three variable dataset. The plots show the distribution of the metrics across the different datasets for each experiment. The centre line is the median, the black dot is the mean, and the white dots represent outliers. Lower is better for SHD and SID, and higher is better for F1.
@@ 5 Related Work
Certain approaches only use information about independences between variables to discover the causal structure, recovering only an MEC, and not the whole DAG (Spirtes et al., 2000, 1995; Chickering, 2002; Huang et al., 2018). Identifying a unique DAG requires additional assumptions, typically involving restrictions on model classes or access to interventional data (Lippe et al., 2021; Ke et al., 2022; Brouillard et al., 2020; Ke et al., 2020). Since we assume access only to observational data, we focus on the former.
Continuous optimisation methods cast causal discovery as a single unconstrained optimisation problem (Zheng et al., 2018). NOTEARS (Zheng et al., 2018) first introduced a regulariser for learning acyclic causal graphs in linear causal models. Follow-up work extended these methods to more flexible models such as neural networks, or graph neural networks (Lachapelle et al., 2019; Yu et al., 2019; Zheng et al., 2020). At the same time, works such as NOBEARS (Lee et al., 2019) and SDCD (Nazaret et al., 2024) have worked on improving the scalability of the acyclic regularisers. However, the above methods rely on restricted models to recover a DAG, which can limit performance when assumptions are not met (Section 2.3). When these models are not restricted, their guarantees no longer hold.
@@
The Bayesian framework has previously been used for causal discovery but has typically focused on computing a posterior over causal structures to quantify uncertainty (Friedman & Koller, 2003; Heckerman et al., 2006). Geiger & Heckerman (2002) use Bayesian linear Gaussian models but construct their model such that graphs in the same Markov equivalence class receive the same score. Similarly, Cundy et al. (2021) also only consider posteriors over linear Gaussian models. DiBS (Lorch et al., 2021) assumes Bayesian ANMs and parametrises the space of graphs using latent variables. Annadani et al. (2023) also use sampling methods to sample from the posterior over ANM causal models, but use a permutation based parametrisation of the graph (Charpentier et al., 2021; Yu et al., 2021).
Previous Bayesian methods have also relied on the restrictive assumptions common in likelihood based methods. In contrast, our work demonstrates that the advantage of Bayesian model selection is the allowance of fewer restrictions compared to likelihood-based methods (Stegle et al., 2010; Dhir et al., 2024; Cooper & Herskovits, 1992). This is strongly related to the idea of using independence of causal mechanisms (and the Kolmogorov complexity of causal models) to distinguish causal models (Janzing & Schölkopf, 2010) (see Dhir et al. (2024, Appendix B)). While previous studies demonstrated this with two variables, we show its feasibility and advantages in the important multi-variable regime.
Figure 2: Box plots of metrics for different graph types: (a) Erdos-Renyi (ER) graphs with four expected edges per variable, (b) Syntren data. Black dot is the average metric. Lower is better for SHD and SID, higher is better for F1.
@@ 6 Experiments
We compare our approach to various baselines that perform multivariate causal discovery to recover a DAG. We test our method on synthetic data generated from our model (Section 6.1), and data not generated from our model (Section 6.2). Then we test our model on a common semi-synthetic benchmark (Section 6.3). Additional experiments and results are in Appendix I. See Section C.6 for full details of the CGP-CDE implementation. Code can be found at: https://github.com/Anish144/ContinuousBMSStructureLearning.git.
It was recently found that causal discovery methods can be sensitive to variance information that is an artefact of synthetic causal datasets (Reisach et al., 2021; Ormaniec et al., 2024). Hence we standardise every variable, including during data generation (see Appendix E).
Baselines: Details for all baselines are in Appendix G. We compare against non-Bayesian counterparts that use acyclicity regularisers, such as NOTEARS (Zheng et al., 2018), and SDCD (Nazaret et al., 2024). SDCD uses the same spectral acyclicity regulariser as the CGP-CDE, the only difference being that the model uses a neural network and a likelihood score to learn the data distribution (Lachapelle et al., 2019). Thus, comparison against this method provides direct evidence for the usefulness of the Bayesian approach. We also compare against methods that first estimate a topological order of the graph and then prune to get the final graph - SCORE (Rolland et al., 2022), NOGAM (Montagna et al., 2023), and CAM (Bühlmann et al., 2014). These all assume ANM. Finally, we compare against the MAP estimate of the Bayesian model DiBS. Other Bayesian methods are sample based and hence do not allow for easy MAP calculation.
Metrics: We use the following metrics to compare the methods. SHD: The structural Hamming distance is the Hamming distance between the predicted adjacency and the ground truth adjacency matrix. SID: The structural interventional distance counts the number of interventional distributions that are incorrect if the predicted graph is used to form the parent adjustment set instead of the ground truth graph (Peters & Bühlmann, 2015). F1: The F1 score is the harmonic mean of the precision and recall, where an edge is considered the positive class.